# VEAttack: Downstream-agnostic Vision Encoder Attack against Large Vision Language Models

### v1

@@ 4.2 Vision encoder attack
Given the VEAttack framework in Section 4.1, we can treat the attacks in adversarial training algorithms proposed by some robust CLIP works mao2022understanding ; schlarmann2024robust as naive solutions to tackle Eq. (4). These vanilla attacks can be formulated as:
$\mathcal{L}_{cls}:\ \tilde{v}=\mathop{\text{arg}\max}_{\|\delta\|\leq\epsilon}-\mathbb{E}_{j}\left[\cos\left(\tilde{z}_{cls},z_{text}^{j}\right)\right]\quad\mathcal{L}_{2}:\ \tilde{v}=\mathop{\text{arg}\max}_{\|\delta\|\leq\epsilon}\|\tilde{z}_{cls}-z_{cls}\|_{2}^{2},$ (5)
where text features $z_{text}^{j}$ is extracted from user instructions. The proposed naive solutions are all classification-centric designed and target at class token embedding $z_{cls}$ within the CLIP vision encoder $f_{CLIP}=\{z_{cls};f_{V}(v)\}$. Unlike classification tasks, LVLMs typically rely on the image token features $z_{v}=f_{V}(v)$ during inference, so the naive attacks need to result in a perturbed image $\tilde{v}$ first and then indirectly affect the visual feature $\tilde{z}_{v}$. This indirect perturbation may limit the effectiveness of attacks on LVLMs, which we will analyze theoretically.
@@
Consider two kinds of attack targets: 1) If the perturbation is introduced to the single class token $z_{cls}$ and propagates through $z_{cls}\xrightarrow{backward}v\xrightarrow{forward}z_{v}$. Let $\Delta z_{cls}(z_{cls})$, $\Delta z_{v}(z_{cls})$, $\Delta z_{m}(z_{cls})$ denote the perturbations on features $z_{cls}$, $z_{v}$ and $z_{m}$ through the first attack, respectively. 2) If the perturbation is directly introduced to the image token features $z_{v}$, let $\Delta z_{m}(z_{v})$ denote the perturbation on the aligned feature $z_{m}$ through the second attack. Assume the same degree of perturbation on $z_{cls}$ and $z_{v}$ during the two attacks, then the ratio of the effect is given by:
@@
Proof. See Appendix B.
From Proposition 2, we infer that perturbations on the class token $z_{cls}$ have a reduced impact on the image token features $z_{v}$, as the ratio $(3+\epsilon_{V})/{\sqrt{n_{v}}}$ is typically smaller than 1 (e.g., $\sqrt{n_{v}}=16$ in CLIP-ViT-L/14). Consequently, this diminished effect propagates to the aligned features $z_{m}$, weakening the adversarial impact on downstream tasks compared to directly targeting $z_{v}$. In Figure 3, we empirically demonstrate that under various perturbation budgets, the perturbations $\|\Delta z_{v}\|_{F}$ and $\|\Delta z_{m}\|_{F}$ induced by attacking $z_{cls}$ are significantly smaller than those from directly attacking $z_{v}$, further validating the superior effectiveness of targeting $z_{v}$. Therefore, we reformulate the objective function from Eq. (4) to directly target $z_{v}$, expressed as $\max_{\|\delta\|\leq\epsilon}\mathcal{L}(f_{V}(v+\delta))$. As illustrated in Figure 4, VEAttack targets the image token features $z_{v}$ of the vision encoder in a downstream-agnostic context. We adopt cosine similarity as the loss metric for $f_{V}(\cdot)$, as $\ell_{2}$ loss may concentrate perturbations on specific dimensions, with limited impact on $z_{m}$ if the alignment layer is less sensitive to those dimensions. In contrast, cosine similarity holistically perturbs the semantic direction of features $z_{v}$, effectively propagating to $z_{m}$. Thus, with cosine similarity as the optimization metric, the generated examples in our VEAttack can be formulated as:
@@ 5.1 Experimental settings
Attack setup. We evaluate the attack performance of our method and baseline methods on pre-trained LVLMs and conduct PGD attack madry2017towards on the vision encoder CLIP-ViT-L/14111https://github.com/openai/CLIP radford2021learning . Following the white-box attack settings schlarmann2023adversarial ; schlarmann2024robust , the perturbation budgets are set as $\epsilon=2/255$ and $\epsilon=4/255$, attack stepsize $\alpha=1/255$, attack steps are $t=100$ iterations. A subset of 500 randomly chosen images is used for adversarial evaluations, with all samples adopted for clean evaluations. The performance degradation ratio after the attack is utilized to assess the attack effect. We conduct all our experiments on a single NVIDIA-A6000 GPU.
Pre-trained models and tasks. Experiments are conducted on diverse LVLMs where LLaVa1.5-7B liu2023visual , LLaVa1.5-13B liu2023visual and OpenFlamingo-9B (OF-9B) awadalla2023openflamingo are evaluated against VEAttack. In addition to them, we adopt miniGPT4 zhu2023minigpt , mPLUG-Owl2 ye2024mplug , and Qwen-VL Qwen-VL for transfer attack of VEAttack. For the image caption task, we evaluate the performance on COCO lin2014microsoft and Flickr30k plummer2015flickr30k datasets. For the VQA task, our experiments are conducted on TextVQA singh2019towards and VQAv2 goyal2017making . All image caption tasks are evaluated using CIDEr vedantam2015cider score, while VQA tasks are measured by VQA accuracy antol2015vqa . Beyond these tasks, we additionally evaluate our VEAttack on a hallucination benchmark called POPE li2023evaluating , and its evaluation metric is F1-score.
Transfer attack setting. The source models contain variants of the CLIP-ViT-L/14 model, which are pre-trained from OpenAI1, TeCoA mao2022understanding and FARE schlarmann2024robust . Both TeCoA and FARE employ adversarial training with a perturbation budget of $\epsilon=4/255$. The target models encompass these three CLIP variants as well as LVLMs from the pre-trained model settings, which utilize vision encoders other than CLIP. Following the setting of transfer attack dong2019evading ; wei2022towards ; chen2023rethinking , we assess transfer attacks across different CLIP models with $\epsilon=8/255$ and transfer attacks between CLIP and non-CLIP vision encoders with $\epsilon=16/255,\alpha=4/255$. All attacks are conducted with $t=100$ iteration steps.
@@ 5.2 VEAttack on LVLMs
We compare VEAttack with baselines in Table 1 across diverse LVLMs and tasks. For the image caption task, VEAttack demonstrates a highly effective attack, reducing the average CIDEr score on COCO by $94.5\%$ under a perturbation of $\epsilon=4/255$. In the VQA task, our proposed VEAttack method also exhibits competitive performance, reducing the average accuracy on the TextVQA dataset by $75.7\%$ under a perturbation of $\epsilon=4/255$. Conversely, other baselines exhibit lower attack effectiveness compared to our method in the vision encoder setting. On the POPE task with F1-score metric, a value below $50\%$ on a balanced dataset indicates poor performance. VEAttack significantly degrades performance under a perturbation of $\epsilon=4/255$, achieving a $38.0\%$ reduction in F1-score and driving it below $50\%$, indicating the VEAttack could increase hallucination in LVLMs.
Table 1: Comparison of VEAttack with baselines across different LVLMs and diverse tasks.
@@
COCO clean 79.7 115.5 119.2 104.8
$\mathcal{L}_{cls}$ mao2022understanding 33.1 21.4 53.7 41.8 56.4 45.5 47.7(-57.1) 36.2(-68.6)
$\mathcal{L}_{2}$ schlarmann2024robust 26.2 11.7 40.3 25.2 40.0 30.0 35.5(-69.3) 22.3(-82.5)
MIX tu2023many 45.9 25.4 67.5 55.4 73.8 60.1 62.4(-42.4) 47.0(-57.8)
VEAttack 7.5 3.7 10.8 7.1 11.2 6.5 9.8(-95.0) 5.8(-99.0)
@@
$\mathcal{L}_{cls}$ mao2022understanding 27.0 18.0 36.8 29.9 40.0 31.4 34.6(-37.0) 26.4(-45.2)
$\mathcal{L}_{2}$ schlarmann2024robust 19.9 11.1 28.5 20.6 30.9 22.5 26.4(-45.2) 18.1(-53.5)
MIX tu2023many 33.7 18.0 42.1 35.9 41.0 32.2 38.9(-32.7) 28.7(-42.9)
VEAttack 8.7 3.2 10.7 6.3 9.1 5.7 9.5(-62.1) 5.1(-66.5)
@@
$\mathcal{L}_{cls}$ mao2022understanding 15.7 11.4 25.2 18.7 24.7 21.2 21.9(-11.4) 17.1(-16.2)
$\mathcal{L}_{2}$ schlarmann2024robust 9.2 7.3 20.1 17.5 17.9 13.0 15.7(-17.6) 12.6(-20.7)
MIX tu2023many 13.4 8.8 24.6 19.1 22.8 19.7 20.3(-13.0) 15.9(-17.4)
VEAttack 12.5 5.7 13.8 10.1 12.4 8.6 12.9(-20.4) 8.1(-25.2)
@@
$\mathcal{L}_{cls}$ mao2022understanding 39.0 38.7 59.7 57.4 61.5 57.1 53.4(-12.8) 51.1(-15.1)
$\mathcal{L}_{2}$ schlarmann2024robust 38.5 34.1 52.9 49.8 54.3 50.4 48.6(-17.6) 44.8(-21.4)
MIX tu2023many 39.8 36.0 59.4 57.9 59.8 58.0 53.0(-13.2) 50.6(-15.6)
VEAttack 34.0 32.8 42.9 38.4 41.5 37.6 39.5(-26.7) 36.3(-29.9)
POPE clean 68.4 84.5 84.1 79.0
$\mathcal{L}_{cls}$ mao2022understanding 66.2 65.1 56.4 55.5 58.5 51.2 60.4(-18.6) 57.3(-21.7)
$\mathcal{L}_{2}$ schlarmann2024robust 65.1 63.6 68.3 61.1 68.9 61.7 67.4(-11.6) 62.1(-16.9)
MIX tu2023many 67.9 67.0 72.0 69.1 74.2 68.2 71.4(-7.6) 68.1(-10.9)
VEAttack 60.6 59.6 47.5 42.8 47.6 44.7 51.9(-27.1) 49.0(-30.0)
@@
Figure 5: t-SNE visualization of the first visual token features across hidden layers in an LLM model for clean and adversarial image inputs.
We visualize the first visual token features of an LLM’s hidden layers using t-SNE in Figure 5 for clean and adversarial samples. To more effectively demonstrate the relative positions of the features in each layer, we denote the positions of the 0th, 16th, and 32nd features. For clean inputs in (a), features show distinct clustering across layers, with progressive spread from Layer 0 to 32. For adversarial inputs in (b), features scatter widely, and the relative relationships between features are more biased. This confirms that attacking the vision encoder induces notable variations in the LLM’s hidden layer representations, despite being in a downstream-agnostic context, validating the effectiveness of the vision encoder attack.
@@
As shown in Table 1, following an identical attack on the vision encoder, the performance degradation in the VQA task is slightly less than that observed in the image caption task. Consequently, we speculate that the VQA task may not pay as much attention to the image tokens as the image caption task. To this end, Figure 6 shows the attention map between the output tokens and the system tokens, image tokens, and instruction tokens within the shallow Layer1, the middle Layer16, and the output Layer32. In the COCO caption task, we observe higher attention scores between output tokens and image tokens, with an average attention value of 0.10 compared to 0.08 in the VQA task. Conversely, in the VQA task, output tokens exhibit stronger attention to instruction tokens, with an average attention value of 0.47, compared to 0.35 in the COCO caption task. This supports our hypothesis that instruction tokens play a more critical role in VQA, while image tokens are more pivotal for image captioning in LVLMs. We hope this observation of differential reliance on image and instruction tokens could guide future researchers in optimizing performance of LVLMs zhang2021vinvl ; xiong2024pyra ; chen2024image .
@@ 5.3 Transfer attack on LVLMs
We first observe transfer attacks using CLIP with varying robustness, as shown in the diagonal performance matrix of Table 2. Employing a more robust CLIP vision encoder significantly enhances LVLM resilience against VEAttack. Specifically, FARE improves average performance by 42.9 over the CLIP of OpenAI, while TeCoA yields a 39.3 improvement. Analyzing the off-diagonal elements of the performance matrix, we observe that adversarial samples crafted against more robust vision encoders exhibit stronger transferability. For instance, attacking FARE reduces the average performance of LVLMs using CLIP as the vision encoder by $58.5\%$, whereas attacking CLIP exhibits limited transferability to other robust vision encoders. Extending our analysis to non-CLIP vision encoders, we find the Möbius band of VEAttack: Strengthening defenses paradoxically fuels more potent attacks, perpetuating a cyclical interplay. Notably, attacking the robust trained vision encoder FARE reduces the performance of miniGPT4 by $90.2\%$ and the performance of Qwen-VL by $25.8\%$. We hope that these findings highlight the critical need for future research to prioritize the security of vision encoders in LVLMs to mitigate adversarial vulnerabilities.
@@
Source Target LLaVa1.5-7B OpenFlamingo-9B miniGPT4 mPLUG-Owl2 Qwen-VL
@@
CLIP radford2021learning 5.2 92.3 103.4 3.2 70.3 79.5 53.3 126.3 131.4
TeCoA mao2022understanding 61.3 44.7 53.9 52.6 42.2 46.8 9.4 101.4 105.4
FARE schlarmann2024robust 47.6 63.6 53.9 33.4 43.1 40.3 9.5 95.5 102.8

### v2

@@ 4.2 Vision encoder attack
Given the VEAttack framework in Section 4.1, we can treat the attacks in adversarial training algorithms proposed by robust CLIP work (Schlarmann et al., 2024) and black-box attack (Zhao et al., 2023) as naive solutions to tackle Eq. (4). These vanilla attacks can be formulated as:
$\text{AttackVLM-ii}:\ \tilde{v}=\mathop{\text{arg}\max}_{\|\delta\|\leq\epsilon}-\cos\left(\tilde{z}_{cls},z_{cls}\right)\quad L_{2}\ \text{Attack}:\ \tilde{v}=\mathop{\text{arg}\max}_{\|\delta\|\leq\epsilon}\|\tilde{z}_{cls}-z_{cls}\|_{2}^{2},$ (5)
The proposed naive solutions are all classification-centric and target at class token embedding $z_{cls}$ within the CLIP vision encoder $f_{CLIP}=\{z_{cls};f_{V}(v)\}$. Unlike classification tasks, LVLMs typically rely on the patch token features $z_{v}=f_{V}(v)$ during inference, so the naive attacks need to result in a perturbed image $\tilde{v}$ first and then indirectly affect the visual feature $\tilde{z}_{v}$. This indirect perturbation may limit the effectiveness of attacks on LVLMs, which we will analyze theoretically.
@@
Consider two kinds of attack targets: 1) If the perturbation is introduced to the single class token $z_{cls}$ and propagates through $z_{cls}\xrightarrow{backward}v\xrightarrow{forward}z_{v}$. Let $\Delta z_{cls}(z_{cls})$, $\Delta z_{v}(z_{cls})$, $\Delta z_{m}(z_{cls})$ denote the perturbations on features $z_{cls}$, $z_{v}$ and $z_{m}$ through the first attack, respectively. 2) If the perturbation is directly introduced to the patch token features $z_{v}$, let $\Delta z_{m}(z_{v})$ denote the perturbation on the aligned feature $z_{m}$ through the second attack. Assume the same degree of perturbation on $z_{cls}$ and $z_{v}$ during the two attacks, then the ratio of the effect is given by:
@@
Proof. See Appendix C.
From Proposition 6, we infer that perturbations on the class token $z_{cls}$ have a reduced impact on the patch token features $z_{v}$, as the ratio $(3+\epsilon_{V})/{\sqrt{n_{v}}}$ is typically smaller than 1 (e.g., $\sqrt{n_{v}}=16$ in CLIP-ViT-L/14). Consequently, this diminished effect propagates to the aligned features $z_{m}$, weakening the adversarial impact on downstream tasks compared to directly targeting $z_{v}$. In Fig. 3, we empirically demonstrate that under various perturbation budgets, the perturbations $\|\Delta z_{v}\|_{F}$ and $\|\Delta z_{m}\|_{F}$ induced by attacking $z_{cls}$ are significantly smaller than those from directly attacking $z_{v}$, further validating the superior effectiveness of targeting $z_{v}$. Therefore, we reformulate the objective function from Eq. (4) to directly target $z_{v}$, expressed as $\max_{\|\delta\|\leq\epsilon}\mathcal{L}(f_{V}(v+\delta))$. As illustrated in Fig. 4, VEAttack targets the patch token features $z_{v}$ of the vision encoder in a downstream-agnostic context. We adopt cosine similarity as the loss metric for $f_{V}(\cdot)$, as $\ell_{2}$ loss may concentrate perturbations on specific dimensions, with limited impact on $z_{m}$ if the alignment layer is less sensitive to those dimensions. In contrast, cosine similarity holistically perturbs the semantic direction of features $z_{v}$, effectively propagating to $z_{m}$. Thus, with cosine similarity as the optimization metric, the generated examples in our VEAttack can be formulated as:
@@ 5.1 Experimental settings
Gray-box attack setup. We evaluate the attack performance of our method and baseline methods on pre-trained LVLMs and conduct PGD attack (Madry et al., 2018) on the vision encoder CLIP-ViT-L/14111https://github.com/openai/CLIP. Following the white-box attack settings (Schlarmann et al., 2024), the perturbation budgets are set as $\epsilon=2/255$ and $\epsilon=4/255$, attack stepsize $\alpha=1/255$, attack steps are $t=100$ iterations. The attack steps of gray-box methods MIX.Attack (Tu et al., 2023) and VT-Attack (Wang et al., 2024c) are $t=1000$. A subset of 500 randomly chosen images is used for adversarial evaluations, with all samples adopted for clean evaluations. The performance degradation ratio after the attack is utilized to assess the attack effect. All experiments were conducted on a NVIDIA-A6000 GPU.
Pre-trained models and tasks. Experiments are conducted on diverse LVLMs where LLaVa1.5-7B (Liu et al., 2023), LLaVa1.5-13B (Liu et al., 2023) and OpenFlamingo-9B (OF-9B) (Awadalla et al., 2023) are evaluated against VEAttack. In addition to them, we adopt miniGPT4 (Zhu et al., 2023), mPLUG-Owl2 (Ye et al., 2024), and Qwen-VL (Bai et al., 2023) for transfer attack of VEAttack. For the image caption task, we evaluate the performance on COCO (Lin et al., 2014) and Flickr30k (Plummer et al., 2015) datasets. For the VQA task, our experiments are conducted on TextVQA (Singh et al., 2019) and VQAv2 (Goyal et al., 2017). All image caption tasks are evaluated using CIDEr (Vedantam et al., 2015) score, while VQA tasks are measured by VQA accuracy (Antol et al., 2015). Beyond these tasks, we additionally evaluate our VEAttack on a hallucination benchmark called POPE (Li et al., 2023b), and its evaluation metric is F1-score.
Transfer attack setting. The source models contain variants of the CLIP-ViT-L/14 model, which are pre-trained from OpenAI1, TeCoA (Mao et al., 2023) and FARE (Schlarmann et al., 2024). Both TeCoA and FARE employ adversarial training with a perturbation budget of $\epsilon=4/255$. The target models encompass these three CLIP variants as well as LVLMs from the pre-trained model settings, which utilize vision encoders other than CLIP. Following the setting of transfer attack (Wei et al., 2022; Chen et al., 2023), we assess transfer attacks across different CLIP models with $\epsilon=8/255$ and transfer attacks between CLIP and non-CLIP vision encoders with $\epsilon=16/255,\alpha=16/(255\times 5)$. All transfer attacks are conducted with $t=100$ iteration steps.
@@ 5.2 VEAttack on LVLMs
We compare VEAttack with gray-box attacks in Table 1 across diverse LVLMs and tasks. For the image caption task, VEAttack demonstrates a highly effective attack, reducing the average CIDEr score on COCO by $94.5\%$ under a perturbation of $\epsilon=4/255$. In the VQA task, our proposed VEAttack method also exhibits competitive performance, reducing the average accuracy on the TextVQA dataset by $75.7\%$ under a perturbation of $\epsilon=4/255$. Conversely, other attacks exhibit lower attack effectiveness compared to our method in the vision encoder setting. On the POPE with F1-score metric, values below $50\%$ on balanced datasets indicate poor performance. VEAttack significantly degrades performance under a perturbation of $\epsilon=4/255$, achieving a $38.0\%$ reduction in F1-score and driving it below $50\%$, indicating the VEAttack could increase hallucination in LVLMs.
Table 1: Comparison of VEAttack with different gary-box attacks across different LVLMs and tasks.
@@
COCO Clean 79.7 115.5 119.2 104.8($\downarrow$0.0%)
MIX.Attack Tu et al. (2023) 45.9 25.4 67.5 55.4 73.8 60.1 62.4($\downarrow$40.5) 47.0($\downarrow$55.1)
VT-Attack Wang et al. (2024c) 38.9 21.6 50.8 12.2 58.2 20.1 49.3($\downarrow$52.9) 18.0($\downarrow$82.8)
AttackVLM-ii Zhao et al. (2023) 24.4 10.3 40.9 25.8 42.7 27.5 36.0($\downarrow$65.6) 21.2($\downarrow$79.8)
VEAttack 7.5 3.7 10.8 7.1 11.2 6.5 9.8($\downarrow$90.6) 5.8($\downarrow$94.5)
@@
MIX.Attack Tu et al. (2023) 33.7 18.0 42.1 35.9 41.0 32.2 38.9($\downarrow$45.7) 28.7($\downarrow$59.9)
VT-Attack Wang et al. (2024c) 27.7 13.8 35.1 12.3 34.0 14.6 32.3($\downarrow$54.9) 13.6($\downarrow$81.0)
AttackVLM-ii Zhao et al. (2023) 18.1 9.9 29.9 19.8 29.7 21.6 25.9($\downarrow$63.8) 17.1($\downarrow$76.1)
VEAttack 8.7 3.2 10.7 6.3 9.1 5.7 9.5($\downarrow$86.7) 5.1($\downarrow$92.9)
@@
MIX.Attack Tu et al. (2023) 13.4 8.8 24.6 19.1 22.8 19.7 20.3($\downarrow$39.0) 15.9($\downarrow$52.3)
VT-Attack Wang et al. (2024c) 15.3 10.5 23.7 10.0 24.9 10.0 21.3($\downarrow$36.0) 10.2($\downarrow$69.4)
AttackVLM-ii Zhao et al. (2023) 12.1 7.6 19.7 11.9 19.8 14.2 17.2($\downarrow$48.3) 11.2($\downarrow$66.4)
VEAttack 12.5 5.7 13.8 10.1 12.4 8.6 12.9($\downarrow$61.3) 8.1($\downarrow$75.7)
@@
MIX.Attack Tu et al. (2023) 39.8 36.0 59.4 57.9 59.8 58.0 53.0($\downarrow$19.9) 50.6($\downarrow$23.6)
VT-Attack Wang et al. (2024c) 38.5 37.0 53.6 21.4 55.2 21.0 49.1($\downarrow$25.8) 26.5($\downarrow$59.9)
AttackVLM-ii Zhao et al. (2023) 37.5 35.8 54.1 49.7 56.2 49.6 49.3($\downarrow$25.5) 45.0($\downarrow$32.0)
VEAttack 34.0 32.8 42.9 38.4 41.5 37.6 39.5($\downarrow$40.3) 36.3($\downarrow$45.2)
POPE Clean 65.7 84.5 84.1 78.1
MIX.Attack Tu et al. (2023) 59.0 53.3 72.0 69.1 74.2 68.2 68.4($\downarrow$12.4) 63.5($\downarrow$18.7)
VT-Attack Wang et al. (2024c) 63.6 63.5 64.0 60.1 65.5 66.4 64.4($\downarrow$17.5) 63.3($\downarrow$18.9)
AttackVLM-ii Zhao et al. (2023) 53.3 48.1 69.0 61.6 63.8 57.8 49.3($\downarrow$36.9) 45.0($\downarrow$42.4)
VEAttack 60.6 59.6 47.5 42.8 47.6 44.7 51.9($\downarrow$33.5) 49.0($\downarrow$37.3)
@@
Figure 5: t-SNE visualization of the features across hidden layers in LLM for clean and adversarial inputs.
We visualize the first visual token features of an LLM’s hidden layers using t-SNE in Fig. 5 for clean and adversarial samples. To more effectively demonstrate the relative positions of the features in each layer, we denote the positions of the 0th, 16th, and 32nd features. For clean inputs in (a), features show distinct clustering across layers, with progressive spread from Layer 0 to 32. For adversarial inputs in (b), features scatter widely, and the relative relationships between features are more biased. This confirms that attacking the vision encoder induces notable variations in the LLM’s hidden layer representations, despite being in a downstream-agnostic context, validating the effectiveness of our vision-encoder-only attack.
@@
As shown in Table 1, following an identical attack on the vision encoder, the performance degradation in the VQA task is slightly less than that observed in the image caption task. Consequently, we speculate that the VQA task may not pay as much attention to the image tokens as the image caption task. To this end, Fig. 6 shows the attention map between the output tokens and the system tokens, image tokens, and instruction tokens within the shallow Layer1, the middle Layer16, and the output Layer32. In the COCO caption task, we observe higher attention scores between output tokens and image tokens, with an average attention value of 0.10 compared to 0.08 in the VQA task. Conversely, in the VQA task, output tokens exhibit stronger attention to instruction tokens, with an average attention value of 0.47, compared to 0.35 in the COCO caption task. This supports our hypothesis that instruction tokens play a more critical role in VQA, while image tokens are more pivotal for image captioning in LVLMs. We hope this observation of differential reliance on image and instruction tokens could guide future researchers in optimizing performance of LVLMs (Zhang et al., 2021; Xiong et al., 2024; Chen et al., 2024).
@@ 5.3 Transfer attack on LVLMs
We first observe transfer attacks using CLIP with varying robustness, as shown in the diagonal performance matrix of Table 2. Employing a more robust CLIP vision encoder significantly enhances LVLM resilience against VEAttack. Specifically, FARE improves average performance by 42.9 over the CLIP of OpenAI, while TeCoA yields a 39.3 improvement. Analyzing the off-diagonal elements of the performance matrix, we observe that adversarial samples crafted against more robust vision encoders exhibit stronger transferability. For instance, attacking FARE reduces the average performance of LVLMs using CLIP as the vision encoder by $58.5\%$, whereas attacking CLIP exhibits limited transferability to other robust vision encoders. Extending our analysis to non-CLIP vision encoders, we find the Möbius band of VEAttack: Strengthening defenses paradoxically fuels more potent attacks, perpetuating a cyclical interplay. Notably, attacking the robust trained vision encoder FARE reduces the performance of miniGPT4 by $90.8\%$ and the performance of mPLUG-Owl2 by $33.5\%$. We hope that these findings highlight the critical need for future research to prioritize the security of vision encoders in LVLMs to mitigate adversarial vulnerabilities.
@@
Source Target LLaVa1.5-7B OpenFlamingo-9B MiniGPT-4 mPLUG-Owl2 Qwen-VL
@@
CLIP Radford et al. (2021) 3.7 92.2 100.4 3.6 70.7 78.1 55.2 124.9 131.2
TeCoA Mao et al. (2023) 22.1 25.3 25.1 17.7 20.8 22.6 11.0 96.4 102.7
FARE Schlarmann et al. (2024) 38.7 54.5 48.6 26.3 42.7 36.4 8.9 88.0 101.4
To provide a mechanistic explanation for the ”Möbius Band” effect (Observation 3), we visualize the dynamics of feature deviation during the attack process. Specifically, we track the difference between the adversarial patch token values and the clean token values across 100 attack steps. We compare three scenarios under a unified metric scale, as shown in Figure 7, which are 1) Attack CLIP, Evaluate on CLIP; 2) Attack FARE, Evaluate on CLIP; 3) Attack CLIP, Evaluate on FARE. When attacking the robust FARE model, the heatmap shows the deepest blue color, appearing early in the attack steps. This indicates that because FARE is resistant to small, high-frequency noise, the optimization process is forced to induce large deviations in the semantic feature space to successfully fool it. When these strong perturbations are transferred to the standard CLIP model, they cause catastrophic feature distortion, explaining the high transfer success rate. In contrast, when attacking CLIP and transferring to FARE, the heatmap remains consistently light. This suggests that the perturbations optimized on CLIP are relatively superficial, which may sufficient to fool CLIP’s decision boundary, but fail to penetrate the robustness of FARE. The robust encoder effectively filters out these specific, lower-magnitude perturbations, resulting in minimal feature deviation and attack failure. In summary, Möbius paradox shows that robust models force the attacker to learn universally destructive features, whereas standard models allow for model-specific fragile features.
Figure 7: Feature deviation heatmaps of tokens across attack steps under different transfer settings.